OpenStudy (anonymous):

what are the first five terms of the sequence given by the formula. An=5n+1 A.1,5,9,13,17 B.6,11,16,21,26 C.5,10,15,20,25 D.5,6,7,8,9 write a recursive definition for the sequence 11,8,5,2 A. a1=11, An=An-1-4 B.a1=11, An=An-1-3 c. a1=2, An=An-1-3 d. a1=2, An=an-1+3

OpenStudy (anonymous):

Really need the help!!

OpenStudy (trojanpoem):

B

OpenStudy (trojanpoem):

Simply , put n = 1 A = 6 , n = 2 A = 11 and so on !

OpenStudy (anonymous):

first term at n=1 A1=5(1)+1=6 second term at n=2 A2=5(2)+1=11

OpenStudy (ilovecake):

B.

OpenStudy (anonymous):

b to what